At highway speeds, a particular automobile is
capable of an acceleration of about 1.6 m/s2. At this
rate, how long does it take to accelerate from 80
km/h to 110 km/h?

Answer by rothauserc(4718)

v = v0 + at, where v is the final velocity, v0 is initial velocity, a is acceleration and t is time, solve for t
:
t = (v - v0) / a
:
note that all variables must be in the same units, we want m/s which is a
:
110 km / hour = (110 * 1000) / (60 * 60) = 30.5556 m/s
:
80 km / hour = (80 * 1000) / (60 * 60) = 22.2222 m/s
:
**********************************************
t = (30.5556 - 22.2222) / 1.6 = 5.2084 seconds
**********************************************
: